OpenStudy (anonymous):

Help me differentiate y^4-xy^2 So far I have 4y^3dy/dx for y^4.... I get stuck on what rule to use on -xy^2 Chain rule?

OpenStudy (aum):

product rule.

OpenStudy (aum):

\[\large (y^4-xy^2)' = 4y^3y' - \{x * 2yy' + y^2 * 1\} = 4y^3y' - 2xyy'- y^2 \\ \]

OpenStudy (aum):

\[ \frac{d}{dx}xy^2 = x * \frac{d}{dx}y^2 + y^2 * \frac{d}{dx}x = x * 2yy' + y^2 * 1 \]

OpenStudy (anonymous):

thank you !